# Modeling Large Nonstationary Spatial Data with the Full-Scale Basis Graphical Lasso

### v1

@@ Abstract
We propose a new approach for the modeling large datasets of nonstationary spatial processes that combines a latent low rank process and a sparse covariance model. The low rank component coefficients are endowed with a flexible graphical Gaussian Markov random field model. The utilization of a low rank and compactly-supported covariance structure combines the full-scale approximation and the basis graphical lasso; we term this new approach the full-scale basis graphical lasso (FSBGL). Estimation employs a graphical lasso-penalized likelihood, which is optimized using a difference-of-convex scheme. We illustrate the proposed approach with a challenging high-resolution simulation dataset of the thermosphere. In a comparison against state-of-the-art spatial models, the FSBGL performs better at capturing salient features of the thermospheric temperature fields, even with limited available training data.
Keywords: graphical lasso, needlets, full-scale approximations
@@ 1 Introduction
The field of spatial statistics has embraced basis function expressions of $Z(x)$ as a useful, interpretable and flexible approach. Endowing the basis functions with random coefficients allows for flexibility in handling different types of process behavior that can also allow for feasible estimation and inference schemes even with very large datasets. Some well-known examples include fixed rank kriging ([frk]), LatticeKrig ([nychka]), the stochastic partial differential equation (SPDE) approach ([lind:11, inla, sigrist]), and the basis graphical lasso ([bgl, mgl]). A comprehensive review of these methods is provided by [annurev:/content/journals/10.1146/annurev-statistics-040120-020733].
In the late 2000s, fixed rank kriging ([frk]) considered a specification of $Z(x)$ as a basis function expansion, with parameters estimated by least squares. Although convenient for large datasets, fixed rank kriging has been criticized for its inability to capture small scale dependence ([steinlowrank]). This work inspired [nychka] in which $Z(x)$ has a very large number of compactly supported functions in a multiresolution expansion which can even exceed the number of observations – the key regularization idea was to enforce a spatial autoregressive (SAR) structure on the coefficient graphical model, thereby reducing the parameter space and affording computational feasibility using sparse matrix methods. This proposal, LatticeKrig, was shown to approximate standard stationary process covariance models under appropriate structuring of the SAR graphical component. A similar concept is applied in [nngp], where a nearest-neighbor structure in space rather than among the coefficients allows the dataset to be described using a directed acyclic graph. By approximating the full density at each point by conditioning on some number of points a reference set (for example, a randomly chosen subset of the spatial locations) instead of the entire dataset, the authors are able to preserve the proper rank of the covariance matrix while making parameter estimation feasible.
In [bgl], the LatticeKrig approach was extended to accommodate nonstationary Gaussian graphical models by imposing a graphical lasso regularization term in a likelihood-estimation framework. Their approach allowed for a nonstationary model that can be efficiently estimated for very large spatial datasets. This approach is termed the basis graphical lasso (BGL).
When the basis expression contains a small number of functions (that is, small compared to the size of the dataset), it is helpful to include an additional stochastic process that represents small-scale behavior ([steinlowrank]). This leads us to consider the representation of the process $Z(x)$ as
@@
where both $Z_{1}$ and $Z_{2}$ are mean zero, spatially correlated processes. The key vision for $Z_{1}$ and $Z_{2}$ is that $Z_{1}$ aims to capture large-scale nonstationary structures via a low-rank basis expression, while $Z_{2}$ is designed for small-scale locally-correlated behavior which we assume to be stationary. Endowing $Z_{2}(x)$ with a compactly supported covariance structure gives rise to the so-called full scale approximation considered in [steinfullscale] and [sangfullscale]. This mixture gives rise to a flexible modeling framework that has a feasible estimation procedure for large datasets. This was expanded upon via a fused Gaussian process model in [ma.kang], where the process $Z_{2}(x)$ is broken up into $M$ disjoint regions described by independent conditional autoregressions that are then fit to the data. This makes fitting the model much more manageable, as instead of fitting to an entire large dataset, which even with a sparse model can be unwieldy, the fitting is done on $M$ smaller datasets.
@@
Further extensions were considered in [katzfuss2013],[katzfuss2017], and [katzfuss2020] with the M-RA, where instead of a single low-rank process the domain is divided up via a recursive domain partitioning scheme or a set of multiresolution tapering functions. This model takes the full-scale approximation of [sangfullscale] and effectively iteratively applies it, so that the full-scale approximation can be seen as the M-RA with one level of resolution ([katzfuss2017]). However, unlike the full-scale approximation, which only rely on a covariance model for the small-scale process and uses empirical orthognal functions (EOFs) derived from the sample covariance for the low-rank portion, this method relies on a covariance model for the entire process, which is then used to determine the basis functions. Additionally, processes on disjoint domains are independent in this construction, so domains must be carefully chosen to respect the assumed covariance length scales.
@@
In this work, we combine the approaches of [bgl] and [sangfullscale] to yield a flexible and interpretable approach to handling large observational datasets of nonstationary spatial processes. Our proposal encapsulates ideas from fixed rank kriging, the full scale approximation, and the basis graphical lasso within one single framework that both allows for flexible nonstationary covariance structures that can also capture small scale correlation structure. For this reason, we call our model the full-scale basis graphical lasso (FSBGL).
@@
To demonstrate the efficacy of this model we assess its ability to represent the statistical structure of lower-thermospheric neutral temperature fields simulated by a large Earth system atomosphere model ([liu.gravity.waves]). The temperature in the lower thermosphere is highly spatially structured and temporally variable due to the influence of atmospheric waves originating from meteorological weather events. Thermospheric gases exert considerable aerodynamic forces on space vehicles and objects, and the spatial structures of lower thermosphere temperature fields considerably affect their reentry trajectory due to their relationship to the neutral mass density ([leonardetal]). The model developed in this study can help engineers, who usually rely on standard atmospheric gas models, account for the effects of spatially structured mass density fields on the expected forces experienced by vehicles during reentry.
@@ 2.1 Statistical Model
for a set of spanning functions $\{\psi_{j}(x)\}_{j=0}^{\infty}\subset\mathcal{L}^{2}(\Omega)$ and where the coefficients $\beta_{j}$ are jointly distributed according to a mean zero multivariate Gaussian graphical model. Specifically, we will let $\beta=(\beta_{0},\beta_{1},\ldots,\beta_{J})^{T}$ where $\beta\sim N(0,Q^{-1})$ with 00 a mean vector of zeros of length $J+1$ and $Q$ the precision matrix. It is important to note that we assume $Q$ is sparse, which will be enforced through regularization in our proposed estimation scheme. Next, let $Z_{2}$ be an isotropic Gaussian process with compactly supported isotropic covariance function $C_{2}(\|h\|)={\rm Cov}(Z_{2}(x+h),Z_{2}(x))$. Finally, let $\varepsilon$ be a mean zero Gaussian white noise process with variance $\tau^{2}$. The major obstacles given this model are in specifying the structure of the graphical model corresponding to $Q$ and estimating parameters of the covariance function $C_{2}$. Synthesizing the assumptions in (1), (2) and (3), our full proposed model for the observation process is
@@ 2.2 Estimation
where $D$ captures the compactly cupported portion of the covariance as well as the white noise, $S$ is the sample covariance matrix of the data, and $\Psi$ is a matrix such that $\Psi_{jk}=\psi_{j}(x_{k})$. The maximum-likelihood estimates of $Q$ and $D$ are the matrices $\tilde{Q}$ and $\tilde{D}$ that satisfy $S=\Psi\tilde{Q}^{-1}\Psi^{T}+\tilde{D}$, however these matrices are difficult to estimate, dense, and often are not uniquely determined by the data.
@@
In the fixed-rank kriging scheme this is handled by setting $Z_{2}=0$, and then $\tau^{2}$ and the covariance of $\{\beta_{j}\}$ is estimated via least-squares ([frk]). If $\Psi$ has a QR decomposition, then [frk] provides closed-form solutions. In contrast, LatticeKrig solves this via parameterization of $Q$ as a block-diagonal set of stationary Gaussian Markov random field precision matrices, specified as spatial autoregressions. In LatticeKrig the basis is fixed to be Wendland Radial Basis Functions and $Q$ is a paraeterized spatial autoregression with independent levels of resolution. The exact computational strategy is outlined in [nychka]. Analogous to fixed rank kriging, within our modeling framework this is captured by setting $Z_{2}(x)$ identically to zero.
@@
One drawback of the LatticeKrig approach is that the stochastic coefficients are restricted to (approximate) stationarity using the spatial autoregressive GMRF. The basis graphical lasso (BGL) relaxes the stationarity assumption, allowing for $Q$ to represent a GMRF structure through a sparsity pattern that is estimated by applying a graphical lasso penalty to the likelihood ([bgl]). This is equivalent to attempting to recover a graphical model for the conditional dependencies among the coefficients, and it means that the whole (potentially dense) coefficient covariance matrix need not be directly estimated. Assuming that $D=\tau^{2}I$, the BGL minimizes the penalized log-likelihood
$L(Q,\tau^{2}|Y)=\log\det\left(\Psi Q^{-1}\Psi^{T}+\tau^{2}I\right)+\\ \text{tr}\left(S\left(\Psi Q^{-1}\Psi^{T}+\tau^{2}I\right)^{-1}\right)+||\Lambda\circ Q||_{1}$ (6)
where $\left|\left|\Lambda\circ Q\right|\right|_{1}=\sum_{i,j}\left|\Lambda_{ij}Q_{ij}\right|$.
@@
In Proposition 1 in [bgl] it was shown that if $\tau^{2}$ is known, minimizing the log-likelihood in (6) is equivalent to minimizing the log-likelihood given by
$\begin{split}L(Q|Y,\tau^{2})&=\log\det\left(Q+\tau^{-2}\Psi^{T}\Psi\right)-\log\det\left(Q\right)\\ &-\textnormal{tr}\left(\tau^{-4}\Psi^{T}S\Psi\left(Q+\tau^{-2}\Psi^{T}\Psi\right)^{-1}\right)+||\Lambda\circ Q||_{1}\end{split}$ (7)
@@
which can be minimized via a difference of convex scheme outlined therein. There, $\tau^{2}$ was calculated with the assumption that $Q=\alpha I$, and then this value was used in the minimization of (7). This model was shown to give superior predictions to other state of the art models such as LatticeKrig and fixed-rank kriging when applied to various climatological datasets. It was also shown to be able to accurately recover the structure of a known precision matrix ([bgl]). However, assuming $D$ to be diagonal is again equivalent to setting $Z_{2}$ to identically zero in our setup.
@@
In [steinfullscale] and [sangfullscale], two similar models are proposed for large spatial datasets. These assume that the field has the form given by (2), where the process $Z(x)=Z_{1}(x)+Z_{2}(x)$ has a low-rank and small-scale portion. These sorts of models have been called full-scale approximations based on terminology from [sangfullscale]. This leads to the covariance structure
$\Psi^{T}\Sigma\Psi+D$ (8)
where $\Psi$ is a matrix of basis functions, $\Sigma$ the covariance matrix of the coefficients, and $D$ the covariance of $Z_{2}$. In [sangfullscale], this was simplified by using empirical orthogonal functions to represent the low rank portion of the model. This keeps $\Sigma$ a diagonal matrix that is estimated simultaneously with $\Psi$.
@@
A key insight in the full-scale approximation is to endow $Z_{2}$ with a compactly supported covariance function, making $D$ sparse. Then the likelihood (5) can be rewritten using the Woodbury formula and properties of the determinant, and parameters of $D$ are estimated from this using MCMC. In [steinfullscale], the basis functions were taken to be Legendre polynomials, and the Cholesky factor of $\Sigma$ was estimated after the parameters of $D$. This was done with a small number of basis functions, keeping estimation simple. As noted in [steinfullscale], estimating a dense Cholesky factorization is $O(n^{3})$, meaning that even though this allows for non-trivial relationships between the basis functions, the estimation procedure is unwieldy for large sets of basis functions without further modification.
@@ 2.2.1 Estimation for the Full-Scale Basis Graphical Lasso (FSBGL)
Both the full-scale approximations and the BGL rely on simplification of the covariance structure, either by keeping $Q$ diagonal via orthonormal basis functions such as in [sangfullscale], keeping the number of basis functions extremely small to allow for estimation of the Cholesky decomposition of the covariance as in [steinfullscale], or by assuming that $Z_{2}(x)$ is identically zero. Our proposal is to combine both methods by allowing for structure in both $Q$ and $Z_{2}$ while still maintaining computational feasibility.
If, instead of letting $D=\tau^{2}I$, we allow $D$ to be an arbitrary positive-definite matrix, then the regularized likelihood is given in the following proposition, a special case of which appeared in [mgl].
@@
$\begin{split}L(Q|Y,D)&=\log\det\left(Q+\Psi^{T}D^{-1}\Psi\right)-\log\det\left(Q\right)\\ &-\textnormal{tr}\left(\Psi^{T}D^{-1}SD^{-1}\Psi\left(Q+\Psi^{T}D^{-1}\Psi\right)^{-1}\right)+||\Lambda\circ Q||_{1}\end{split}$ (9)
The proof is presented in Appendix A, but it is a straightforward generalization of the proof presented in [bgl]. This gives the field covariance the form of a low-rank representation plus a high-resolution component represented by $D$, which is similar to the full-scale approximations discussed in [steinfullscale] and [sangfullscale]. The matrix solves in (9) can be made efficient by endowing $Z_{2}$ with a compactly supported covariance function, imposing sparsity in $D$. For this reason, we call this model the full-scale basis graphical lasso (FSBGL).
As suggested by [bgl], we can write the likelihood as a sum of convex and concave parts, and we can solve the optimization for $Q$ by a Difference of Convex (DC) scheme. The convex part is given by $-\log\det(Q)+||\Lambda\circ Q||_{1}$ and the concave part by $\log\det\left(Q+\Psi^{T}D^{-1}\Psi\right)-\textnormal{tr}\left(\Psi^{T}D^{-1}SD^{-1}\Psi\left(Q+\Psi^{T}D^{-1}\Psi\right)^{-1}\right)$, and so the DC scheme for minimization is given by
$\hat{Q}_{j+1}|\hat{Q}_{j},D$ $=\underset{Q>0}{\text{argmin}}\left[-\log\det(Q)+\text{tr}(\nabla g(\hat{Q}_{j})Q)+||\Lambda\circ Q||_{1}\right]$
@@
The derivation is provided in Appendix B. Successive calculations of $\hat{Q}_{j+1}$ are performed using QUIC ([JMLR:v15:hsieh14a]). We say that the method has converged when, for a given $\delta$, $\frac{\left|\left|\hat{Q}_{j+1}-\hat{Q}_{j}\right|\right|_{Fro}}{\left|\left|\hat{Q}_{j}\right|\right|_{Fro}}<\delta$. In this paper, we set $\delta=0.02$.
To fit the model, we need to first produce an estimate of $D$ to be used in the minimization in (10). To do this, we adapted the algorithm used in [bgl] to fit the nugget variance $\tau^{2}$. We make the assumption that $Q=\alpha I$, and then jointly fit $Q$ and $D$. We prefer to estimate this way because there is not a clear path to estimating a full model for $Q$ and $D$. Moreover, based on initial experiments, re-estimating $D$ after estimating $Q$ was not shown to lead to significant improvement, as also seen in [bgl].
@@
which can be derived as an intermediate step in the derivation of (9). This can be fit by any standard optimization method. Once $\hat{p}$ is known, the precision matrix $Q$ is fit via the optimization problem in Equation (10), with the penalty matrix given by
@@
Since we fit the model for a range of $\lambda$, we need a model selection criterion to determine the optimal value of $\lambda$ for each model. In order to avoid overfitting, we choose to use the conditional Akaike information criterion in the form ([hatmatrixAIC])

### v2

@@ Abstract
We propose a new approach for the modeling large datasets of nonstationary spatial processes that combines a latent low rank process and a sparse covariance model. The low rank component coefficients are endowed with a flexible graphical Gaussian Markov random field model. The utilization of a low rank and compactly-supported covariance structure combines the full-scale approximation and the basis graphical lasso; we term this new approach the full-scale basis graphical lasso (FSBGL). Estimation employs a graphical lasso-penalized likelihood, which is optimized using a difference-of-convex scheme. We illustrate the proposed approach on synthetic fields as well as with a challenging high-resolution simulation dataset of the thermosphere. In a comparison against state-of-the-art spatial models, the FSBGL performs better at capturing salient features of the thermospheric temperature fields, even with limited available training data.
keywords:
graphical lasso, needlets, full-scale approximations
journal: Spatial Statistics
[…]
@@ 1 Introduction
The field of spatial statistics has embraced basis function expressions of $Z(x)$ as a useful, interpretable and flexible approach. Endowing the basis functions with random coefficients allows for flexibility in handling different types of process behavior that can also allow for feasible estimation and inference schemes even with very large datasets. Some well-known examples include fixed rank kriging ([4]), LatticeKrig ([31]), the stochastic partial differential equation (SPDE) approach ([23, 2, 34]), and the basis graphical lasso ([20, 21]). A comprehensive review of these methods is provided by [5].
In the late 2000s, fixed rank kriging ([4]) considered a specification of $Z(x)$ as a basis function expansion, with parameters estimated by least squares. Although convenient for large datasets, fixed rank kriging has been criticized for its inability to capture small scale dependence ([37]). This work inspired [31] in which $Z(x)$ has a very large number of compactly supported functions in a multiresolution expansion which can even exceed the number of observations – the key regularization idea was to enforce a spatial autoregressive (SAR) structure on the coefficient graphical model, thereby reducing the parameter space and affording computational feasibility using sparse matrix methods. This proposal, LatticeKrig, was shown to approximate standard stationary process covariance models under appropriate structuring of the SAR graphical component. A similar concept is applied in [6], where a nearest-neighbor structure in space rather than among the coefficients allows the dataset to be described using a directed acyclic graph. By approximating the full density at each point by conditioning on some number of points a reference set (for example, a randomly chosen subset of the spatial locations) instead of the entire dataset, the authors are able to preserve the proper rank of the covariance matrix while making parameter estimation feasible. Another similar concept is applied in [15] to the problem of modeling intercropping data with asymmetric effects across species, with the graphical model describing relationships between crop species rather than over space.
In [20], the LatticeKrig approach was extended to accommodate nonstationary Gaussian graphical models by imposing a graphical lasso regularization term in a likelihood-estimation framework. Their approach allowed for a nonstationary model that can be efficiently estimated for very large spatial datasets. This approach is termed the basis graphical lasso (BGL).
When the basis expression contains a small number of functions (that is, small compared to the size of the dataset), it is helpful to include an additional stochastic process that represents small-scale behavior ([37]). This leads us to consider the representation of the process $Z(x)$ as
@@
where both $Z_{1}$ and $Z_{2}$ are mean zero, spatially correlated processes. The key vision for $Z_{1}$ and $Z_{2}$ is that $Z_{1}$ aims to capture large-scale nonstationary structures via a low-rank basis expression, while $Z_{2}$ is designed for small-scale locally-correlated behavior which we assume to be stationary. Endowing $Z_{2}(x)$ with a compactly supported covariance structure gives rise to the so-called full scale approximation considered in [36] and [33]. This mixture gives rise to a flexible modeling framework that has a feasible estimation procedure for large datasets. This was expanded upon via a fused Gaussian process model in [26], where the process $Z_{2}(x)$ is broken up into $M$ disjoint regions described by independent conditional autoregressions that are then fit to the data. This makes fitting the model much more manageable, as instead of fitting to an entire large dataset, which even with a sparse model can be unwieldy, the fitting is done on $M$ smaller datasets.
@@
Further extensions were considered in [17],[18], and [19] with the M-RA, where instead of a single low-rank process the domain is divided up via a recursive domain partitioning scheme or a set of multiresolution tapering functions. This model takes the full-scale approximation of [33] and effectively iteratively applies it, so that the full-scale approximation can be seen as the M-RA with one level of resolution ([18]). However, unlike the full-scale approximation, which only rely on a covariance model for the small-scale process and uses empirical orthognal functions (EOFs) derived from the sample covariance for the low-rank portion, this method relies on a covariance model for the entire process, which is then used to determine the basis functions. Additionally, processes on disjoint domains are independent in this construction, so domains must be carefully chosen to respect the assumed covariance length scales.
@@
In this work, we combine the approaches of [20] and [33] to yield a flexible and interpretable approach to handling large observational datasets of nonstationary spatial processes. Our proposal encapsulates ideas from fixed rank kriging, the full scale approximation, and the basis graphical lasso within one single framework that both allows for flexible nonstationary covariance structures that can also capture small scale correlation structure. For this reason, we call our model the full-scale basis graphical lasso (FSBGL).
@@
To demonstrate the efficacy of this model we assess its ability to represent the statistical structure of lower-thermospheric neutral temperature fields simulated by a large Earth system atomosphere model ([25]). The temperature in the lower thermosphere is highly spatially structured and temporally variable due to the influence of atmospheric waves originating from meteorological weather events. Thermospheric gases exert considerable aerodynamic forces on space vehicles and objects, and the spatial structures of lower thermosphere temperature fields considerably affect their reentry trajectory due to their relationship to the neutral mass density ([22]). The model developed in this study can help engineers, who usually rely on standard atmospheric gas models, account for the effects of spatially structured mass density fields on the expected forces experienced by vehicles during reentry.
@@ 2.1 Statistical Model
for a set of spanning functions $\{\psi_{j}(x)\}_{j=0}^{\infty}\subset\mathcal{L}^{2}(\Omega)$ and where the coefficients $\beta_{j}$ are jointly distributed according to a mean zero multivariate Gaussian graphical model. Specifically, we will let $\beta=(\beta_{0},\beta_{1},\ldots,\beta_{J})^{T}$ where $\beta\sim N(0,Q^{-1})$ with 0 a mean vector of zeros of length $J+1$ and $Q$ the precision matrix. It is important to note that we assume $Q$ is sparse, which will be enforced through regularization in our proposed estimation scheme. Next, let $Z_{2}$ be an isotropic Gaussian process with compactly supported isotropic covariance function $C_{2}(\|h\|)={\rm Cov}(Z_{2}(x+h),Z_{2}(x))$. Finally, let $\varepsilon$ be a mean zero Gaussian white noise process with variance $\tau^{2}$. The major obstacles given this model are in specifying the structure of the graphical model corresponding to $Q$ and estimating parameters of the covariance function $C_{2}$. Synthesizing the assumptions in (1), (2) and (3), our full proposed model for the observation process is
@@ 2.2 Estimation
where $D$ captures the compactly cupported portion of the covariance as well as the white noise, $S$ is the sample covariance matrix of the data, and $\Psi$ is a matrix such that $\Psi_{ij}=\psi_{j}(x_{i})$. The maximum-likelihood estimates of $Q$ and $D$ are the matrices $\tilde{Q}$ and $\tilde{D}$ that satisfy $S=\Psi\tilde{Q}^{-1}\Psi^{T}+\tilde{D}$, however these matrices are difficult to estimate, dense, and often are not uniquely determined by the data.
@@
In the fixed-rank kriging scheme this is handled by setting $Z_{2}=0$, and then $\tau^{2}$ and the covariance of $\{\beta_{j}\}$ is estimated via least-squares ([4]). If $\Psi$ has a QR decomposition, then [4] provides closed-form solutions. In contrast, LatticeKrig solves this via parameterization of $Q$ as a block-diagonal set of stationary Gaussian Markov random field precision matrices, specified as spatial autoregressions. In LatticeKrig the basis is fixed to be Wendland Radial Basis Functions and $Q$ is a paraeterized spatial autoregression with independent levels of resolution. The exact computational strategy is outlined in [31]. Analogous to fixed rank kriging, within our modeling framework this is captured by setting $Z_{2}(x)$ identically to zero.
@@
One drawback of the LatticeKrig approach is that the stochastic coefficients are restricted to (approximate) stationarity using the spatial autoregressive GMRF. The basis graphical lasso (BGL) relaxes the stationarity assumption, allowing for $Q$ to represent a GMRF structure through a sparsity pattern that is estimated by applying a graphical lasso penalty to the likelihood ([20]). This is equivalent to attempting to recover a graphical model for the conditional dependencies among the coefficients, and it means that the whole (potentially dense) coefficient covariance matrix need not be directly estimated. Assuming that $D=\tau^{2}I$, the BGL minimizes the penalized log-likelihood
[…]
@@
In Proposition 1 in [20] it was shown that if $\tau^{2}$ is known, minimizing the log-likelihood in (6) is equivalent to minimizing the log-likelihood given by
[…]
@@
which can be minimized via a difference of convex scheme outlined therein. There, $\tau^{2}$ was calculated with the assumption that $Q=\alpha I$, and then this value was used in the minimization of (7). This model was shown to give superior predictions to other state of the art models such as LatticeKrig and fixed-rank kriging when applied to various climatological datasets. It was also shown to be able to accurately recover the structure of a known precision matrix ([20]). However, assuming $D$ to be diagonal is again equivalent to setting $Z_{2}$ to identically zero in our setup.
@@
In [36] and [33], two similar models are proposed for large spatial datasets. These assume that the field has the form given by (2), where the process $Z(x)=Z_{1}(x)+Z_{2}(x)$ has a low-rank and small-scale portion. These sorts of models have been called full-scale approximations based on terminology from [33]. This leads to the covariance structure
$\Psi\Sigma\Psi^{T}+D$ (8)
where $\Psi$ is a matrix of basis functions, $\Sigma$ the covariance matrix of the coefficients, and $D$ the covariance of $Z_{2}$.
@@
A key insight in the full-scale approximation is to endow $Z_{2}$ with a compactly supported covariance function, making $D$ sparse. Then the likelihood (5) can be rewritten using the Woodbury formula and properties of the determinant, and parameters of $D$ are estimated from this using MCMC. In [36], the basis functions were taken to be Legendre polynomials, and the Cholesky factor of $\Sigma$ was estimated after the parameters of $D$. This was done with a small number of basis functions, keeping estimation simple. As noted in [36], estimating a dense Cholesky factorization is $O(n^{3})$, meaning that even though this allows for non-trivial relationships between the basis functions, the estimation procedure is unwieldy for large sets of basis functions without further modification. In [33], this was simplified by using empirical orthogonal functions to represent the low rank portion of the model. This keeps $\Sigma$ a diagonal matrix that is estimated simultaneously with $\Psi$. Unfortunately neither of these approaches work well for a large number of basis functions as the covariance matrix $\Sigma$ would require some regularization, or severe limiting assumptions (e.g., being diagonal as in [33]). In this work we relax these restrictions by incorporating lasso-like regularization on the inverse covariance matrix of the stochastic coefficients.
@@ 2.2.1 Estimation for the Full-Scale Basis Graphical Lasso (FSBGL)
Both the full-scale approximations and the BGL rely on simplification of the covariance structure, either by keeping $Q$ diagonal via orthonormal basis functions such as in [33], keeping the number of basis functions extremely small to allow for estimation of the Cholesky decomposition of the covariance as in [36], or by assuming that $Z_{2}(x)$ is identically zero. Our proposal is to combine both methods by allowing for structure in both $Q$ and $Z_{2}$ while still maintaining computational feasibility.
If, instead of letting $D=\tau^{2}I$, we allow $D$ to be an arbitrary positive-definite matrix, then the regularized likelihood is given in the following proposition, a special case of which appeared in [21].
@@
[…]
The proof is presented in Appendix A, but it is a straightforward generalization of the proof presented in [20]. This gives the field covariance the form of a low-rank representation plus a high-resolution component represented by $D$, which is similar to the full-scale approximations discussed in [36] and [33]. The matrix solves in (9) can be made efficient by endowing $Z_{2}$ with a compactly supported covariance function, imposing sparsity in $D$. For this reason, we call this model the full-scale basis graphical lasso (FSBGL).
As in [20], the main motivation for using Eq. (9) rather than Eq. (5) to fit the model is computational complexity. If we are to use Eq. (5) to fit the model,at each step we must compute the Cholesky decomposition of the matrix $\Psi Q^{-1}\Psi^{T}+D$, which is an $O(n^{3})$ operation that is not helped by the fact that this matrix will, in general, be dense. Then once the Cholesky factor $L$ is calculated, the log-determinant is calculated in $O(n)$ and the trace can be calculated as $\frac{1}{n}\|L^{-1}Y\|_{F}^{2}$, which is $O(n^{2}m+n^{2})$. Additionally, the calculation of $\Psi Q^{-1}\Psi$ is $O(n^{2}J+J^{2}n+J^{3})$, leading to overall complexity that is $O(n^{3})$.
In our preferred implementation, Eq. (9), we are able to leverage the sparsity of $D$ to improve the situation. First, calculating the Cholesky decomposition of $D=L_{D}L_{D}^{T}$ is an $O(n^{3})$ operation naively, however since it is known it only needs to be done once, and the sparsity of $D$ will accelerate both this decomposition and the resulting applications of $L_{D}$ greatly compared to the previous case where all involved matrices were dense. In the event that $D$ is a banded matrix, for example, this can be done in $O(k^{2}n)$ time, where $k$ is the number of non-zero diagonals. If $D$ is block-diagonal this will carry into the Cholesky decomposition, increasing the speed of those calculations as well. Then, the factors $\Psi^{T}D^{-1}\Psi$ and $\Psi^{T}D^{-1}SD^{-1}\Psi$ can be precomputed as well, and in fact since $tr(A^{T}A)=\|A\|_{F}^{2}$ the trace term can be made even simpler to compute, requiring $O(J^{3}n^{2})$ operations per step of the algorithm and only requiring that $\Psi^{T}D^{-1}Y$ be precomputed rather than $\Psi^{T}D^{-1}SD^{-1}\Psi$. The remaining operations are at worst $O(J^{3})$, leading to significant computational improvements.
[…]
@@
[…]
To fit the model, we need to first produce an estimate of $D$ to be used in the minimization in (10). To do this, we adapted the algorithm used in [20] to fit the nugget variance $\tau^{2}$. We make the assumption that $Q=\alpha I$, and then jointly fit $\alpha$ and $D$. We prefer to estimate this way because there is not a clear path to estimating a full model for $Q$ and $D$ simultaneously. Once we have determined $D$, we can then use Eq. (10) to get the final estimate of $Q$. Based on initial experiments, re-estimating $D$ after estimating a full $Q$ was not shown to lead to significant improvement, as also seen in [20].
@@
which can be derived as an intermediate step in the derivation of (9). This can be fit by any standard optimization method, although care should be taken to make sure that operations on $D$ do not become prohibitively expensive. One option for this would be to parameterize $D$ by a Markov Random Field or other method that allows direct parameterization of the inverse, another might be to give $D$ block-diagonal structure or directly parameterize the Cholesky decomposition, as done in [29], although if one is not careful the latter could greatly increase the number of parameters to estimate in return. We also recommend applying a derivative-free optimization method to avoid the need to calculate finite-difference approximations to the gradient, which in general will not have a closed form, thus avoiding extra likelihood evaluations. In the event that $n$ is small and there are a small number of parameters, this is less of a concern.
Once $\hat{p}$ is known, the precision matrix $Q$ is fit via the optimization problem in Equation (10), with the penalty matrix given by
@@
Since we fit the model for a range of $\lambda$, we need a model selection criterion to determine the optimal value of $\lambda$ for each model. In order to avoid overfitting, we choose to use the conditional Akaike information criterion in the form ([32])